\begin{thebibliography}{99}
%\bibitem{CRS} D. Cvetkovi\'c, P. Rowlinson and S. Simi\'c, An intrdoction to the theory of graph spectra, London Mathematical Society Student Text %Books,
%{\bf 75}, Cambridge University Press, New York, 2010.
%\bibitem{S} D. Stevanovi\'c, Research problems from the Aveiro workshop on graph spectra, {\em Linear Algebra and its Applications}, {\bf 423} (2007)
%172-181.

\bibitem{ajo} A. Abdollahi, Sh. Janbaz and M. R. Oboudi, Distance between spectra of graphs, Linear Algebra Appl., 466 (2015), 401--408.

\bibitem{ao} A. Abdollahi and M. R. Oboudi, Cospectrality of graphs, Linear Algebra Appl., 451 (2014), 169--181.

\bibitem{ahk} M. Afkhami, M. Hassankhani and K. Khashyarmanesh, Distance between the spectra of graphs with respect to normalized Laplacian spectra, Georgian Math. J., 26 (2017), No. 2, 227--234. 

\bibitem{CY} D. Cao and H. Yuan, Graphs characterized by the second eigenvalue, J. Graph Theory,
17 (1993), No. 3, 325--331.

\bibitem{ccgh} G. Caporossi, D. Cvetkovi\'c, I. Gutman and P. Hansen, Variable neighbourhood search for extermal graphs. 2. Finding graphs with extremal energy, J. Chem. Inf. Comput. Sci., 39 (1999), 984--996.

\bibitem{clss} H. Choi, H. Lee, Y. Shen and Y. Shi, Comparing large-scale graphs based on quantum probability theory, Appl. Math. Comput., 358 (2019), 1--15.

%\bibitem{crs} D. Cvetkovi\'c, P. Rowlinson, S. Simi\'c, An Introduction to the Theory of Graph spectra, London Mathematical Society Student Texts, 75. Cambridge University Press, Cambridge, 2010.

\bibitem{CDS} D. M. Cvetkovi\'c, M. Doob and H. Sachs, Spectra of Graphs, Theory and Application, Academic Press, Inc., New York, 1979.

%Cvetković, Dragoš M.; Doob, Michael; Sachs, Horst. Spectra of graphs. Theory and applications. Third edition. Johann Ambrosius Barth, Heidelberg, 1995. 
\bibitem{ds} K. Ch. Das and S. Sun, Distance between the normalized Laplacian spectra of two graphs, Linear Algebra Appl., 530 (2017), 305--321.

\bibitem{lld} H. Lin, D. Li and K. Ch. Das, Distance between distance spectra of graphs, Linear Multilinear Algebra, 65 (2017), No. 12, 2538--2550.

\bibitem{Godsil} C. Godsil and G. Royle, Algebraic Graph Theory, Springer, New York, 2001.

%Godsil, Chris; Royle, Gordon. Algebraic graph theory. Graduate Texts in Mathematics, 207. Springer-Verlag, New York, 2001.

\bibitem{ghl} J. Gu, B. Hua and S. Liu, Spectral distances on graphs, Discrete Appl. Math., 190/191 (2015), 56--74.

\bibitem{ha} M. Hakimi-Nezhaad and A. R. Ashrafi, Laplacian and normalized Laplacian spectral distances of graphs, Southeast Asian Bull. Math., 37 (2013), No. 5, 731--744.

\bibitem{j} I. Jovanovi\'c, Some results on spectral distances of graphs, Rev. Un. Mat. Argentina, 56 (2015), No. 2, 95--117.

\bibitem{js} I. Jovanovi\'c and Z. Stani\'c, Spectral distances of graphs, Linear Algebra Appl., 436 (2012), No. 5, 1425--1435.

\bibitem{jzs} I. Jovanovi\'c and Z. Stani\'c, Spectral distances of graphs based on their different matrix representations, Filomat, 28 (2014), No. 4, 723--734. 

\bibitem{o} M. R. Oboudi, Cospectrality of complete bipartite graphs, Linear Multilinear Algebra, 64 (2016), No. 12, 2491--2497.

\bibitem{p} M. Petrovi\'c, On graphs whose second largest eigenvalue does not exceed $\sqrt{2}-1$, Univ. Beograd. Publ. Elektrotehn. Fak. Ser. Mat., 4 (1993), 70--75.

\bibitem{S} J. H. Smith, Symmetry and multiple eigenvalues of graphs, Glasnik Mat., Ser. III, 12 (1977) No. 1, 3--8.

%Smith, J. H. Symmetry and multiple eigenvalues of graphs. Glasnik Mat. Ser. III 12(32) (1977), no. 1, 3--8.

\bibitem{s} D. Stevanovi\'c, Research problems from the Aveiro Workshop on Graph Spectra, Linear Algebra Appl., 423 (2007), No. 1, 172--181.

\bibitem{wz} R.C. Wilson and P. Zhu, A study of graph spectra for comparing graphs and trees,
Pattern Recogn., 41 (2008), 2833--2841.

%A study of graph spectra for comparing graphs and trees, Pattern Recogn., 41 (9) (2008), pp. 2833-2841

\bibitem{zp} K. Zając and J. Piersa, Eigenvalue Spectra of Functional Networks in fMRI Data and Artificial Models. In: Rutkowski L., Korytkowski M., Scherer R., Tadeusiewicz R., Zadeh L.A., Zurada J.M. (eds) Artificial Intelligence and Soft Computing. ICAISC 2013. Lecture Notes in Computer Science, vol 7894. Springer, Berlin, Heidelberg, 2013. 

%Zając, K., & Piersa, J. (2013). Eigenvalue spectra of functional networks in fMRI data and artificial models. ICAISC 2013, PT I. Lecture Notes in Artificial Intelligence, 7894, 205–214.

\end{thebibliography}

\documentclass{amsart}
\usepackage[top=1 in,bottom=1.25 in,left=2cm, right=2cm]{geometry}
\usepackage{amsmath, amsthm, amscd, amsfonts, amssymb, color}
\usepackage[dvips]{graphicx}
\usepackage{tikz}
\usepackage{float}
\usepackage{perpage}
\usepackage[T1]{fontenc}
\newtheorem{thm}{Theorem}[section]

\newtheorem{lem}[thm]{Lemma}
\newtheorem{prop}[thm]{Proposition}
\theoremstyle{definition}

\numberwithin{equation}{section}

\begin{document}

\title[$\ell^1$-Cospectrality of graphs]{$\ell^1$-Cospectrality of graphs}%
\author[A. Abdollahi]{Alireza Abdollahi}%
\address{Department of Mathematics, University of Isfahan, Isfahan 81746-73441, Iran; and School of Mathematics, Institute for Research in Fundamental Sciences (IPM), P.O. Box 19395-5746, Tehran, Iran}%
\email{a.abdollahi@math.ui.ac.ir}%
\author[N. Zakeri]{Niloufar Zakeri}
\address{Department of Mathematics, University of Isfahan, Isfahan 81746-73441, Iran}%
\email{zakeri@sci.ui.ac.ir}%
\thanks{}%
\subjclass[2010]{05C50; 05C31}%
\keywords{Spectra of graphs; Cospectrality of graphs; Adjacency matrix of a graph; $\ell^1$-norm}%

\begin{abstract}
The following problem  has been proposed in  [Research problems from the Aveiro workshop on graph spectra, {\em Linear Algebra and its Applications}, {\bf 423} (2007)
172-181.]:\\
(Problem AWGS.4) Let $G_n$ and $G'_n$ be two nonisomorphic graphs on $n$ vertices with spectra
$$\lambda_1 \geq \lambda_2 \geq \cdots \geq \lambda_n \;\;\;\text{and}\;\;\; \lambda'_1 \geq \lambda'_2 \geq \cdots \geq \lambda'_n,$$
respectively. Define the distance between the spectra of $G_n$ and $G'_n$ as
$$\lambda(G_n,G'_n) =\sum_{i=1}^n (\lambda_i-\lambda'_i)^2 \;\;\; \big(\text{or use}\; \sum_{i=1}^n|\lambda_i-\lambda'_i|\big).$$
Define the cospectrality of $G_n$ by
$$\text{cs}(G_n) = \min\{\lambda(G_n,G'_n) \;:\; G'_n \;\;\text{not isomorphic to} \; G_n\}.$$
{\bf Problem A.} Investigate $\text{cs}(G_n)$ for special classes of graphs.\\

In this paper we study Problem A for certain graphs with respect to the $\ell^1$-norm, i.e. $\sigma(G_n,G'_n)=\sum_{i=1}^n|\lambda_i-\lambda'_i|$. We find $\text{cs}(K_n)$, $\text{cs}(nK_1)$, $\text{cs}(K_2+(n-2)K_1)$ ($n\geq 2$), $\text{cs}(K_{n,n})$ and $\text{cs}(K_{n,n+1})$, where $K_n, nK_1, K_2+(n-2)K_1, K_{n,m} $ denote the complete graph on $n$ vertices,  the null graph on $n$ vertices, the disjoint union of the $K_2$ with $n-2$ isolated vertices  ($n\geq 2$), and the complete bipartite graph with parts of sizes $n$ and $m$, respectively.
\end{abstract}
\maketitle
\section{\bf Introduction and Results}
Throughout the paper all graphs are simple, that is finite and undirected without loops and multiple edges.\\
Richard Brualdi proposed in  \cite{s} the following problem:\\
(Problem AWGS.4) Let $G_n$ and $G'_n$ be two nonisomorphic graphs on $n$ vertices with spectra
$$\lambda_1 \geq \lambda_2 \geq \cdots \geq \lambda_n \;\;\;\text{and}\;\;\; \lambda'_1 \geq \lambda'_2 \geq \cdots \geq \lambda'_n,$$
respectively. Define the distance between the spectra of $G_n$ and $G'_n$ as
$$\lambda(G_n,G'_n) =\sum_{i=1}^n (\lambda_i-\lambda'_i)^2 \;\;\; \big(\text{or use}\; \sum_{i=1}^n|\lambda_i-\lambda'_i|\big).$$
Define the cospectrality of $G_n$ by
$$\text{cs}(G_n) = \min\{\lambda(G_n,G'_n) \;:\; G'_n \;\;\text{not isomorphic to} \; G_n\}.$$
Let $$\text{cs}_n = \max\{\text{cs}(G_n) \;:\; G_n \;\;\text{a graph on}\; n \;\text{vertices}\}.$$
This function measures how far apart the spectrum of a graph with $n$ vertices can be from the
spectrum of any other graph with $n$ vertices.\\

\noindent {\bf Problem A.} Investigate $\text{cs}(G_n)$ for special classes of graphs.\\
{\bf Problem B.} Find a good upper bound on $\text{cs}_n$.\\

See, for example \cite{clss,wz,zp} some applications of spectral distances of graphs. 
In \cite{ajo}, Problem B has completely been answered. 
It is of course possible to study Problems A or B for other matrix representations of graphs, such as Laplacian, normalized Laplacian, signless Laplacian and distance matrices, see e.g., \cite{ahk,lld,ds,ghl,ha,j,jzs}. In the current paper, we only study Problem A with respect to the adjacency matrix of graphs.

In \cite{ao,o}, Problem A is studied and  cospectralities of  classes of complete graphs and complete bipartite  graphs with respect to Euclidean norm (the $\ell^2$-norm) are computed.

In \cite{js}, spectral distance between certain graphs is studied with respect to the $\ell^1$-norm i.e. $\sigma(G_n,G'_n)=\sum_{i=1}^n|\lambda_i-\lambda'_i|$. 

In this paper we study Problem A for some graphs with respect to the $\ell^1$-norm. \\

Let us first introduce some notations.\\

For a graph $G$, $V(G)$ and $E(G)$ denote the vertex set and edge set of $G$, respectively; By the  order and size of $G$ we mean the number of vertices and the number of edges of $G$, respectively; Denote by $\overline{G}$ the complement of $G$. Let $G$ be a graph with vertex set $\{v_1,\dots,v_n\}$. The adjacency matrix of $G$ is an $n\times n$ matrix $A(G)=[a_{ij}]$ such that $a_{ij}=1$ if $v_i$ and $v_j$ are adjacent, and $a_{ij}=0$ otherwise. By the eigenvalues of $G$, we mean those of its adjacency matrix. We denote by $Spec(G)$ the multiset of the eigenvalues of the graph $G$. For two graphs $G$ and $H$ with disjoint vertex sets, $G+H$ denotes the graph with the vertex set $V(G) \cup V(H)$ and the edge set $E(G) \cup E(H)$, i.e. the disjoint union of two graphs $G$ and $H$. The complete product (join) $G \nabla H$ of graphs $G$ and $H$ is the graph obtained from $G+H$ by joining every vertex of $G$ with every vertex of $H$. In particular, $nG$ denotes $\underbrace{G+\cdots+G}_n$ and $\nabla_nG$ denotes $\underbrace{G\nabla G \nabla \dots \nabla G}_n$.

For positive integers $n_1, \dots,n_\ell$, $K_{n_1,\dots,n_\ell}$ denotes the complete multipartite graph with $\ell$ parts of sizes $n_1,\dots,n_\ell$.
Let $K_n$ denote the complete graph on $n$ vertices, $nK_1=\overline{K_n}$ denote the null graph on $n$ vertices and $P_n$ denote the path with $n$ vertices.\\

   
Here we find the cospectralities with respect to the $\ell^1$-norm of those graphs  which have been already found their ones with respect to the $\ell^2$-norm in \cite{ao}. In the following, we give the table of cospectralities of these graphs with respect to the $\ell^1$-norm and $\ell^2$-norm. In the third column (fifth column, resp.)  all graphs $H$ with $\text{cs}(G)=\sigma(G,H)$ ($\text{cs}(G)=\lambda(G,H)$, resp.) are given for the graph $G$ in the first column.  \\

\begin{small}
\begin{tabular}{|c|c|c|c|c|}\hline
Graph & $\ell^1$-norm & $H$ & $\ell^2$-norm & $H$\\ \hline
$nK_1$& 2& $K_2+(n-2)K_1$& 2 & $K_2+(n-2)K_1$\\
$K_2+(n-2)K_1$& $2(\sqrt{2}-1)$&$P_3+(n-3)K_1$&$2(\sqrt{2}-1)^2$&$P_3+(n-3)K_1$\\
$K_n$&2&$K_n\setminus e$ or $K_{n-1}+K_1$ &$n^2+n-n\sqrt{n^2+2n-7}-2$& $K_n\setminus e$\\
$K_{n,n}$&$ 2(n-\sqrt{n^2-1})$&$K_{n-1,n+1}$&$2(n-\sqrt{n^2-1})^2$&$K_{n-1,n+1}$\\
$K_{n,n+1}$&$2(\sqrt{n^2+n}-\sqrt{n^2+n-2})$&$K_{n-1,n+2}$&$2(\sqrt{n^2+n}-\sqrt{n^2+n-2})^2$&$K_{n-1,n+2}$\\\hline
\end{tabular}\\
\end{small}

 Our main results are as follows.

\begin{thm}\label{empty}
 For every integer $n\geq2$, $\text{cs}(nK_1)=2$. Moreover, $\text{cs}(nK_1)=\sigma(nK_1,H)$ for some graph $H$ if and only if $H\cong K_2+(n-2)K_1$.
\end{thm}

\begin{thm}\label{oneedge}
$\text{cs}(K_2)=\sigma(K_2,2K_1)=2$ and for every integer $n\geq3$, $\text{cs}(K_2+(n-2)K_1)=2(\sqrt{2}-1)$. Moreover,  $\text{cs}(K_2+(n-2)K_1)=\sigma(K_2+(n-2)K_1,H)$ for some graph $H$ if and only if $H\cong P_3+(n-3)K_1$.
\end{thm}

\begin{thm}\label{kn}
Let $n\geq 2$ be an integer. Then $\text{cs}(K_n)=2$. Moreover, $\text{cs}(K_n)=\sigma(K_n,H)$ for some graph $H$ if and only if $H\cong K_{n-1}+K_1$ or $H\cong K_n\setminus e$ for any edge $e$, where $K_n\setminus e$ is the graph obtaining from $K_n$ by deletion one edge $e$.
\end{thm}

\begin{thm}\label{knnn}
For every integer $n\geq 2$, $\text{cs}(K_{n,n})=2(n-\sqrt{n^2-1})$. Moreover, $\text{cs}(K_{n,n})=\sigma(K_{n,n},H)$ for some graph $H$ if and only if
$H\cong K_{n-1,n+1}$.
\end{thm}

\begin{thm}\label{knn+1}
For every integer $n\geq 2$, $\text{cs}(K_{n,n+1})=2(\sqrt{n^2+n}-\sqrt{n^2+n-2})$. Moreover, $\text{cs}(K_{n,n+1})=\sigma(K_{n,n+1},H)$ for some graph $H$ if and only if
$H\cong K_{n-1,n+2}$.
\end{thm}

\section{\bf $\ell^1$-Cospectrality  of graphs with at most one edge}
In this section we will determine the cospectrality  of
 the graphs with at most one edge. Let $G$ be a simple graph of order $n$ and $\lambda_1\geq\cdots\geq \lambda_n $ be the eigenvalues of $G$. Recall that the energy of $G$ is defined as $E(G)=\displaystyle\sum_{i=1}^{n} |\lambda_i|$. 
We need the following Theorem in the sequel.\\

\begin{thm}[See \cite{ccgh}]\label{eng} 
Let $G$ be a graph with m edges. Then\\
 $$E(G)\geq 2\sqrt{m},$$ \\
with equality if and only if $G$ is a complete bipartite graph plus arbitrarily isolated vertices.
\end{thm}
 
\noindent{\bf Proof of Theorem \ref{empty}.} Let $H$ be a simple graph of order $n$ and size $m$. Clearly, we have $\sigma(nK_1,H)=E(H)$. Since $H$ is not isomorphic to $nK_1$, by Theorem \ref{eng}, the minimum value of $E(H)$ is 2 and it happens whenever $m=1$. This shows that $H\cong K_2+(n-2)K_1$. $\hfill \Box$ \\

\noindent{\bf Proof of Theorem \ref{oneedge}.} It is easy to see that $\text{cs}(K_2)=\sigma(K_2,2K_1)=2$. Suppose $H$ is a simple graph of order $n$ and size $m$ with eigenvalues $\lambda_1\geq \cdots \geq \lambda_n$. For every integer $n\geq 3$, 
it follows from Theorem \ref{eng} and triangular inequality that 
\begin{eqnarray*}
\sigma(K_2+(n-2)K_1,H)&=&\mid \lambda_1 - 1\mid + \displaystyle\sum_{i=2}^{n-1} \mid \lambda_i \mid + \mid \lambda_n+1 \mid \\
 &=& E(H)-\mid\lambda_1\mid-\mid\lambda_n\mid+\mid\lambda_1-1\mid+\mid\lambda_n+1\mid \\
 &\geq & 2\sqrt{m}-\mid\lambda_1\mid-\mid\lambda_n\mid+|\lambda_1|-1+|\lambda_n|-1\\
 &\geq& 2(\sqrt{m}-1),
\end{eqnarray*}
Note that if $m=0$, then $\sigma(K_2+(n-2)K_1,H)=2$. Now assume that $m>0$. 
Since $H$ is not isomorphic to $K_2+(n-2)K_1$, the minimum value of $\sigma(K_2+(n-2)K_1,H)$ is $2(\sqrt{2}-1)$ and it happens whenever $m=2$. The graphs with two edges are $2K_2+(n-4)K_1$ and $P_3+(n-3)K_1$ and  
$$\sigma(K_2+(n-2)K_1, 2K_2+(n-4)K_1)=2,$$
$$\sigma(K_2+(n-2)K_1, P_3+(n-3)K_1)=2(\sqrt{2}-1).$$
It shows that $H\cong P_3+(n-3)K_1$. This completes the proof. $\hfill \Box$
 
\section{\bf $\ell^1$-Cospectrality of the complete graph}

In this section we show that for every integer $n\geq 2$, the minimum value of $\sigma(K_n,H)$ happens whenever $H\cong K_n\setminus e$ or $K_{n-1}+K_1$, and $\text{cs}(K_n)=2$. We need the following results.

\begin{thm}[\cite{js}, part (i) of Theorem 3.4]\label{jov}
Let $G$ be an arbitrary graph, and let $n^*$ denote the number of its eigenvalues which are greater than or equal to $-1$. Then the following holds:
$$\sigma(K_n,G)=2\big(n^{*}-1+\displaystyle\sum_{i=2}^{n^{*}}\lambda_i(G)\big).$$
\end{thm}

\begin{lem} [See \cite{ao}]\label{speckne}
For every integer $n\geq2$ and every arbitrary edge $e$ of $K_n$,
$$Spec(K_n\setminus e)=\Big\{\frac{n-3+\sqrt{n^2+2n-7}}{2},0,\underbrace{-1,\ldots,-1}_{n-3},\frac{n-3-\sqrt{n^2+2n-7}}{2}\Big\}.$$
\end{lem}

\begin{lem} [See \cite{ao}]\label{lemS}
Let $G$ be a graph with eigenvalues $\lambda_1\geq \lambda_2\geq \lambda_3\geq \cdots \geq \lambda_n$.
 The graph $G$ is isomorphic to one the following graphs if and only if $\lambda_1>0$, $\lambda_2\leq 0$ and $\lambda_3<0$.
\begin{enumerate}
\item $G\cong K_n$,
\item $G\cong K_1+K_{n-1}$, \item $G\cong K_{2,1,\dots,1}=
K_n\setminus e$ for an edge $e$ of $K_n$.
\end{enumerate}
\end{lem}

\noindent{\bf Proof of Theorem \ref{kn}.} The result easily  follows for $n=2$. So we may assume that $n\geq 3$. It is easy to see that $\sigma(K_n,K_{n-1}+K_1)=2$. By Theorem \ref{jov}, Lemma \ref{speckne} and the fact the sum of eigenvalues of a simple graph is zero, we obtain
\begin{eqnarray*}
\sigma(K_n,K_n\setminus e)&=&2\big(n-2+\displaystyle\sum_{i=2}^{n-1}\lambda_i(K_n\setminus e)\big)\\
&=&2 \big(n-2-\lambda_1(K_n\setminus e)-\lambda_n(K_n\setminus e)\big)\\
&=&2.
\end{eqnarray*}
Let $H$ be a simple graph of order $n$ and let $n^*$ denote the number of eigenvalues of $H$ which are greater than or equal to $-1$. To complete the proof, it is sufficient to show that if $\sigma(K_n,H)\leq2$ then $H$ is isomorphic to one of the following graphs: $K_n$, $K_{n-1}+K_1$ or $K_n\setminus e$. By Theorem \ref{jov},
$$\sigma(K_n,H)=2\big(n^*-1+\displaystyle\sum_{i=2}^{n^*}\lambda_i(H)\big).$$
If $\sigma(K_n,H)\leq 2$, it follows that 
\begin{equation}\label{eq1}
n^*+\displaystyle\sum_{i=2}^{n^*}\lambda_i(H)\leq 2.
\end{equation}
Let ${n^*}_{\geq 0}$ and ${n^*}_{<0}$ denote the cardinality of the following sets $A$ and $B$, respectively,
$$A=\{ \lambda_i(H) \mid 0\leq \lambda_i(H) \},$$
$$B=\{ \lambda_i(H) \mid -1\leq \lambda_i(H)<0 \}.$$
By Inequality \ref{eq1}, 
$${n^*}_{\geq 0}+{n^*}_{<0}+\displaystyle\sum_{\lambda_i(H)\in A \; \text{and}\; i>1}\lambda_i(H)+\displaystyle\sum_{\lambda_i(H)\in B}\lambda_i(H)\leq 2.$$
Since ${-n^*}_{<0}\leq\displaystyle\sum_{\lambda_i(H)\in B}\lambda_i(H)$, 
\begin{equation}\label{eq2}
{n^*}_{\geq 0}+\displaystyle\sum_{\lambda_i(H)\in A}\lambda_i(H)\leq 2.
\end{equation}
Since $\sigma(K_n,nK_1)=2n-2$ and $n\geq 3$, we can assume that $H$ has at least one edge. It implies that $\lambda_1(H)> 0$ and so ${n^*}_{\geq 0}\geq 1$. By Inequality \ref{eq2}, we have the following cases:\\

\noindent {\bf Case 1.} \; ${n^*}_{\geq 0}=1$. Then both of $\lambda_2(H)$ and $\lambda_3(H)$ are negative.\\

\noindent {\bf Case 2.} \; ${n^*}_{\geq 0}=2$. Then $\lambda_2(H)=0$ and $\lambda_3(H)<0$.\\

Therefore by Lemma \ref{lemS}, $H$ is isomorphic to one of the following graphs: $K_n$, $K_{n-1}+K_1$ or $K_n\setminus e$. This completes the proof. $\hfill \Box$

\section{\bf $\ell^1$-Cospectrality of complete bipartite graphs}
We need the following results to prove  Theorem \ref{knnn}.

\begin{thm}[Theorem 9.1.1 of \cite{Godsil}]\label{interlacing} Let $G$ be a graph of order $n$ and $H$ be an induced subgraph of $G$ with order $m$. Suppose that $\lambda_1(G)\geq\cdots\geq\lambda_n(G)$ and $\lambda_1(H)\geq\cdots\geq\lambda_m(H)$ are the eigenvalues of $G$ and $H$, respectively. Then for every $i$, $1\leq i\leq m$, $\lambda_i(G)\geq\lambda_i(H)\geq\lambda_{n-m+i}(G)$.
\end{thm}

\begin{thm}[Theorem 1 of \cite{CY}] \label{eigenvalue}
Let $G$ be a simple graph of order $n$ without isolated vertices. If $\lambda_2(G)$ is the second largest eigenvalue of $G$, then
\begin{enumerate}
\item $\lambda_2(G)=-1$ iff $G$ is a complete graph with at least two vertices.
\item $\lambda_2(G)=0$ iff $G$ is a complete $k$-partite graph with $2\leq k\leq n - 1$.
\item There exists no graph $G$ such that $-1 \leq \lambda_2(G) \leq 0$.
\end{enumerate}
\end{thm}

\begin{thm}[See \cite{S}, and also Theorem 6.7 of \cite{CDS}] \label{Sm} A graph has exactly one positive eigenvalue if and only if its non-isolated vertices form a complete multipartite graph.
\end{thm}

\begin{thm}[Theorem 2 of \cite{CY}] \label{1/3} Let $G$ be a graph of order $n$ without isolated vertices. Then $0<\lambda_2(G)<\frac{1}{3}$
if and only if $G\cong (K_1+K_2)\nabla\overline{K_{n-3}}$, where
$\lambda_2(G)$ is the second largest eigenvalue of $G$.
\end{thm}

\begin{prop}[Proposition 4.1 of \cite{ao}]\label{kmn}
Let $m$ and $n$ be positive integers. Then $\text{cs}(K_{m,n})>0$ if and only if  the minimum of $x+y$ for all positive integers  $x,y$ such that $xy=mn$
is  attained on $\{x,y\}=\{m,n\}$.
\end{prop}

By Proposition \ref{kmn}, if $n>0$, then $\text{cs}(K_{n,n})>0$. Now we compute  $\text{cs}(K_{n,n})$.\\ 
The method that we use below for the proof is similar to that of Theorem 1.4 of \cite{ao}.\\

\noindent{\bf Proof of Theorem \ref{knnn}.}
Since for all positive integers $p$ and $q$ 
$$ Spec(K_{p,q})=\{\sqrt{pq},\underbrace{0,\ldots,0}_{p+q-2},-\sqrt{pq}\},$$  
it implies that $\sigma(K_{n,n},K_{n-1,n+1})=2(n-\sqrt{n^2-1})$.
By direct computing, the result follows for $n=2$ and $n=3$. Now we can assume that $n\geq 4$. Suppose, for a contradiction, that there exists a graph $H$ of order $2n$ such that $H$ is not isomorphic to either $K_{n,n}$ or  $K_{n-1,n+1}$, and also $\sigma(K_{n,n},H)\leq 2(n-\sqrt{n^2-1})$. Let $\lambda_1\geq\cdots\geq\lambda_{2n}$ be the eigenvalues of $H$.
 Since for $n\geq 4$, $2(n-\sqrt{n^2-1})<\frac{1}{3}$, it follows that $|\lambda_2|<\frac{1}{3}$.
By Theorem \ref{eigenvalue}, we have $0\leq\lambda_2<\frac{1}{3}$. Now it remains to investigate the following cases:\\

\noindent {\bf Case 1.} \; Assume that $\lambda_2=0$. If $\lambda_1=0$, then $H\cong \overline{K_{2n}}$ and so
$\sigma(K_{n,n},H)=2n\geq8$, a contradiction. Hence we can suppose that $\lambda_1>0$. By Theorem \ref{Sm},
there exist some positive integers $k$ and $n_1,\dots,n_k$ and an
integer $t\geq0$ such that $H\cong \overline{K_t}+K_{n_1,\ldots,n_k}$. If
$k=1$, then $H\cong \overline{K_t}+K_{2n-t}$ such that $1\leq t\leq 2n-2$. We have
$$\sigma(K_{n,n},H)=\mid n-t-1\mid+ 3n-t-3.$$
One can to see that $\sigma(K_{n,n},H)>2(n-\sqrt{n^2-1})$, a contradiction. If $k=2$, then
$H\cong \overline{K_t}+K_{p,q}$ for some $p$ and $q$ such that $p+q=2n-t$.
In this case we have $\sigma(K_{n,n},H)=2\mid n-\sqrt{pq}\mid$. It is
not difficult to see that if $\{p,q\}\not=\{n,n\}$ and
$\{p,q\}\not=\{n-1,n+1\}$, then
$2\mid n-\sqrt{pq}\mid >2(n-\sqrt{n^2-1})$. Therefore $k\geq 3$. If
$n_1=\cdots=n_k=1$, then $H\cong\overline{K_t} + K_{2n-t}$ and so
$\sigma(K_{n,n},H)> 2(n-\sqrt{n^2-1})$, a contradiction. Now we
may assume that $n_i\geq 2$, for some $1\leq i \leq k$. So $H$ has $K_{1,1,2}$ as an induced subgraph. Since $Spec(K_{1,1,2})=\{2.56155, 0, -1, -1.56155\}$ and
$\lambda_3(K_{1,1,2})=-1$, by 
Theorem \ref{interlacing}, we have $\mid\lambda_{2n-1}\mid\geq 1$ and so
$\sigma(K_{n,n},H)\geq 1$, a contradiction. \\

\noindent {\bf Case 2.} \; Suppose that $0<\lambda_2<\frac{1}{3}$. By Theorem \ref{1/3}, there exists an integer $t\geq0$ such that $H\cong
\overline{K_t}+(K_1+K_2)\nabla\overline{K_{2n-t-3}}$. Let $2n-t-3=1$. Since $Spec((K_1+K_2)\nabla K_1)=\{2.17009, .31111, -1, -1.48119\}$, it is not hard to see that $\sigma(K_{n,n},H)>2(n-\sqrt{n^2-1})$, a
contradiction. So we can assume that $2n-t-3>1$. Then $H$ has $K_{1,1,2}$ as an induced subgraph and the rest is similar to the previous part.\\
This completes the proof. $\hfill\Box$ \\

We need the following results to proof Theorem \ref{knn+1}.

\begin{thm}[See {\cite{p}}]\label{petro}  
Let $G$ be a graph without isolated vertices and let $\lambda_2(G)$ be the second largest eigenvalue of $G$. Then $0<\lambda_2(G)\leq\sqrt{2}-1$ if and only if one of the following holds:
\begin{enumerate}
\item $G\cong (\nabla_{t}(K_1+K_2))\nabla K_{n_1,\ldots,n_m},$
\item $G\cong (K_1+K_{r,s})\nabla\overline{K_q},$
\item $G\cong (K_1+K_{r,s})\nabla K_{p,q}.$
\end{enumerate}
\end{thm}
Now we prove an ``$\ell^1$-version'' of  Lemma 2.7 of \cite{o}.
\begin{lem}\label{knm}
Let $m$ and $n$ be two positive integers and $G$ be a graph of order $n + m$. Suppose that there are no positive integers $r, s$ and a non-negative integer $t$ such that
$G\cong K_{r,s}+tK_1$. If $\lambda_2(G)\leq \sqrt{2}-1$, then $\sigma(G,K_{m,n})\geq 1$.
\end{lem}
\begin{proof}
By Theorem \ref{eigenvalue}, we can consider the following cases:\\

\noindent {\bf Case 1.} \; $\lambda_2(G)=-1$. Thus $G\cong K_{n+m}$. It is not hard to see that $\sigma(G,K_{m,n})\geq 1$.\\

\noindent {\bf Case 2.} \; $\lambda_2(G)=0$. If $\lambda_1(G)=0$, then $G\cong \overline{K_{m+n}}$. Therefore 
 $\sigma(G,K_{m,n})=2\sqrt{mn}\geq 2$. Now we suppose that $\lambda_1(G) > 0$. By Theorem \ref{Sm}, there are some positive integers $k, n_1,\ldots, n_k$ and a non-negative integer t such that $G\cong K_{n_1,\ldots,n_k}+tK_1$. If $k=1$, then $G\cong K_{n+m-t}+tK_1$. We have 
$$\sigma(G,K_{m,n})=|m+n-t-1-\sqrt{mn}|+(m+n-t-2)+|1-\sqrt{mn}|,$$
Since $\lambda_2(G)=0$ and $\lambda_1(G)> 0$, $1\leq t\leq m+n-2$ and $m,n\geq 2$. So $\sigma(G,K_{m,n})\geq 1$. If $k=2$, then $G\cong K_{n_1,n_2}+tK_1$, a contradiction. Let $k\geq 3$. If $n_1=\cdots=n_k=1$, then $G\cong K_{n+m-t}+tK_1$. Therefore $\sigma(G,K_{m,n})\geq 1$. Now we can assume that $n_i\geq 2$, for some $1\leq i \leq k$. So $G$ has $K_{1,1,2}$ as an induced subgraph. Since $\lambda_3(K_{1,1,2})=-1$, by Theorem \ref{interlacing}, we have $\mid\lambda_{n+m-1}(G)\mid\geq 1$ and so $\sigma(G,K_{m,n})\geq 1$. \\

\noindent {\bf Case 3.} \; $0<\lambda_2(G)\leq \sqrt{2}-1$. By Theorem \ref{petro}, $G$ is isomorphic to one the following graphs: $(\nabla_h(K_1+K_2))\nabla K_{n_1,\ldots,n_k}+tK_1$, $(K_1+K_{r,s})\nabla\overline{K_q}+tK_1$ or $(K_1+K_{r,s})\nabla K_{p,q}+tK_1$, where $k, r, s, p, q$ and $n_1,\dots,n_k$ are some positive integers
and $t, h$ are two non-negative integers. Let $G\cong (\nabla_h(K_1+K_2))\nabla K_{n_1,\ldots,n_k}$. If $h=0$, then $G\cong K_{n_1,\ldots,n_k}$, that is considered in the previous part. If $h\geq 1$, then $(K_1+K_2)\nabla K_1$ is an induced sungraph of $G$. Since $\lambda_3((K_1+K_2)\nabla K_1)=-1$, by Theorem \ref{interlacing}, $\mid\lambda_{n+m-1}\mid\geq 1$ and so $\sigma(G,K_{m,n})\geq 1$. Also $(K_1+K_2)\nabla K_1$ is an induced subgraph for other cases and the result follows by Theorem \ref{interlacing}.
\end{proof}
\noindent{\bf Proof of Theorem \ref{knn+1}.}
We have $\sigma(K_{n,n+1},K_{n-1,n+2})=2(\sqrt{n^2+n}-\sqrt{n^2+n-2})$. 
Suppose, for a contradiction, that $H$ is a graph not isomorphic to both of $K_{n,n+1}$, $K_{n-1,n+2}$ and $\text{cs}(K_{n,n+1})=\sigma(K_{n,n+1},H)$. By direct computing of cospectralities of all graphs of orders at most $10$, one finds that the order of $H$ is at least $11$. We may assume that $n=5$. Note that $\sigma(K_{5,6},K_{4,7})=2(\sqrt{30}-\sqrt{28})$. 
If there are positive integers $r$, $s$ and a non-negative integer $t$ such that $H\cong K_{r,s}+tK_1$ and $r+s+t = 11$, then $\sigma(K_{5,6},H)=2|\sqrt{30}-\sqrt{rs}|> 2(\sqrt{30}-\sqrt{28})$.  So we can assume that there are no positive integers $r$, $s$ and a non-negative integer $t$ such that $H\cong K_{r,s}+tK_1$ and $r+s+t= 11$. Since $\sigma(K_{5,6},H)\leq 2(\sqrt{30}-\sqrt{28})<\sqrt{2}-1$, $|\lambda_2(H)|<\sqrt{2}-1$. By Lemma \ref{kmn}, $\sigma(K_{5,6},H)\geq 1$, a contradiction. We conclude that the result holds for $n=5$. Now we may assume that $n\geq 6$ and by similar arguments given in Theorem \ref{knnn}, the result follows. $\hfill\Box$ 

\section*{{\bf Acknowledgments}}

The research of the first author was in part supported by a grant from School of Mathematics, Institute for Research in Fundamental Sciences(IPM).


\end{document}